924 Massachusetts Avenue, �2 Arlington, MA 02174

May 1, 1998

Mitotix Inc. One Kendall Square Bld. 600 Cambridge, MA 02139

Dear Sir or Madam,

I am interested in opportunities within your company as a part- or fulltime senior research assistant or laboratory technologist. I have looked at the job opportunities listing on your web page and would like to be considered for job� ASP98, Associate Scientist in protein puri�cation and production in particular. I have eight years of experience in the �elds of cellular and molecular biology, protein chemistry, and enzyme kinetics. I have recently set up and run an enzyme kinetics laboratory.

I am enclosing a copy of my resume for your perusal, and would be interested in discussing any relevant job openings. Please contact me at (781) 648-0367 or via email at cjsmith@mit.edu. Thank you for your time; I look forward to hearing from you.

Sincerely,

May 1, 1998

NeoGenesis Inc. Human Resources Ref. HN-1 Attn: Biology 840 Memorial Drive Cambridge, MA 02139

Dear Sir or Madam,

I am interested in opportunities within your company as a part- or full-time senior research assistant or laboratory technologist. I have eight years of experience in the �elds of cellular and molecular biology, protein chemistry, and enzyme kinetics. I have recently set up and run an enzyme kinetics laboratory.

I am enclosing a copy of my resume for your perusal, and would be interested in discussing any relevant job openings. Please contact me at (781) 648-0367 or via email at cjsmith@mit.edu. Thank you for your time; I look forward to hearing from you.

Sincerely,

May 1, 1998

The Genetics Institute Human Resources Department 87 Cambridge Park Drive Cambridge, MA 02139

Dear Sir or Madam,

I am interested in opportunities within your company as a part- or full-time senior research assistant or laboratory technologist. I have looked at your job opportunities listed on your web page and I am interested in job� 09-0000025. I have eight years of experience in the �elds of cellular and molecular biology, protein chemistry, and enzyme kinetics. I have recently set up and run an enzyme kinetics laboratory.

I am enclosing a copy of my resume for your perusal, and would be interested in discussing any relevant job openings. Please contact me at (781) 648-0367 or via email at cjsmith@mit.edu. Thank you for your time; I look forward to hearing from you.

Sincerely,